Find each percent increase or decrease. Round each percent to the nearest tenth, if necessary 110 to 40
Answer by nerdybill(7384)

Find each percent increase or decrease. Round each percent to the nearest tenth, if necessary 110 to 40
.
difference: 110-40=70
percent: 70/110*100 = 0.636*100= 63.6%
.
answer: 63.6% DECREASE